OpenStudy (anonymous):

Find the derivative of the function. h(t) = (t4 - 1)^5 (t3 + 1)^8 5(t^4 - 1)^4 (4 * t^3) 8(t^3+1)^7(3*t^2) h'(t) = (t^4 -1)^5 (8(t^3+1)^7(3*t^2))+(t^3+1)^8(5(t^4-1)^4(4 *t^3) Now I am confused with the problem from this point on how to get the answer, the video lecture skips exactly which steps to factor to get the right answer.

OpenStudy (loser66):

\(h'(t) = \color {red}{((t^4-1)^5)' }*(t^3+1)^8+(t^4-1)^5*\color{blue}{((t^3+1)^8)'}\)

OpenStudy (loser66):

\(h'(t) = \color {red}{(5(t^4-1)^4)(4t^3) }*(t^3+1)^8+(t^4-1)^5*\color{blue}{(8(t^3+1)^7(3t^2)}\)

OpenStudy (loser66):

factor out : \(h'(t) = 2t^2(t^4-1)^4(t^3+1)^7[10t(t^3+1)+12(t^4-1)]\)

OpenStudy (anonymous):

The part I am confused with your second answer to your third answer which is the solution. I am confused how it goes to 2t^2 in front and how you get 10t and 12 to the back of the problem.

OpenStudy (loser66):

from 2, the first term: number: 5*4 =20, right? the second term: number: 8*3 = 24, instead of factor 4 out, I just take 2 out, you should take out 2 more , hehehe. my mistake

OpenStudy (loser66):

\(h'(t) = \color {red}{(5(t^4-1)^4)(4t^3) }*(t^3+1)^8\\~~~~~~~~~~~~~~+(t^4-1)^5*\color{blue}{(8(t^3+1)^7(3t^2)}\) rearrange: \(h'(t) = \color {red}{(20(t^4-1)^4)(t^3) }*(t^3+1)^8\\+~~~~~~~~~~24(t^4-1)^5*\color{blue}{((t^2)(t^3+1)^7}\)

OpenStudy (loser66):

Factor the smallest term out to get my third line

OpenStudy (anonymous):

Okay thanks, I finally see what you did in the problem. thanks.